# Motion Planning Diffusion: Learning and Adapting Robot Motion Planning with Diffusion Models

### v1

@@ Abstract
The performance of optimization-based robot motion planning algorithms is highly dependent on the initial solutions, commonly obtained by running a sampling-based planner to obtain a collision-free path. However, these methods can be slow in high-dimensional and complex scenes and produce non-smooth solutions. Given previously solved path-planning problems, it is highly desirable to learn their distribution and use it as a prior for new similar problems. Several works propose utilizing this prior to bootstrap the motion planning problem, either by sampling initial solutions from it, or using its distribution in a maximum-a-posterior formulation for trajectory optimization. In this work, we introduce Motion Planning Diffusion (MPD), an algorithm that learns trajectory distribution priors with diffusion models. These generative models have shown increasing success in encoding multimodal data and have desirable properties for gradient-based motion planning, such as cost guidance. Given a motion planning problem, we construct a cost function and sample from the posterior distribution using the learned prior combined with the cost function gradients during the denoising process. Instead of learning the prior on all trajectory waypoints, we propose learning a lower-dimensional representation of a trajectory using linear motion primitives, particularly B-spline curves. This parametrization guarantees that the generated trajectory is smooth, can be interpolated at higher frequencies, and needs fewer parameters than a dense waypoint representation. We demonstrate the results of our method ranging from simple $2$D to more complex tasks using a $7$-dof robot arm manipulator. In addition to learning from simulated data, we also use human demonstrations on a real-world pick-and-place task. The experiments show that MPD achieves higher success rates compared to an uninformed prior, and a baseline that first samples from the diffusion prior and then optimizes the cost, while maintaining diversity in the generated trajectories.
https://sites.google.com/view/motionplanningdiffusion
@@ I Introduction
In this article, we present Motion Planning Diffusion (MPD) – a method that combines learning and adaptation of robot trajectories using diffusion models. Our contributions are:
  1. 1.
    We learn trajectory distribution priors using diffusion in a compact B-spline parametrization, using data of path planning solutions or real-world human demonstrations;
  2. 2.
    For controlled trajectory generation, we condition our model on a start joint position and goal configurations (desired joint position or an end-effector pose);
  3. 3.
    Given a cost function encoding a motion planning task, we use planning-as-inference and leverage cost guidance in diffusion models to obtain diverse and collision-free trajectories;
  4. 4.
    Our results show that MPD obtains more valid and diverse trajectories than the baselines in various motion planning problems, from simple $2$D up to complex environments with a $7$-dof robot.
@@ II-A Path and motion planning
On the other hand, optimization-based motion planners search for a collision-free movement by locally optimizing a trajectory either via gradient descent or stochastic optimization, aiming for smoothness while satisfying other objective constraints. These methods can be seen as performing sampling-based planners’ search and smoothening steps in one step. Gradient-based methods include CHOMP [2], TrajOpt [4] and GPMP/GPMP2 [3]. Stochastic optimization methods based on path integral formulations [32] include STOMP [33], Stochastic GPMP [34] and MGPTO [35]. Other works, such as MPOT [36], solve motion planning by optimal transport.
@@ II-C Diffusion models in robotics and motion planning
In recent years, several works have explored score-based and diffusion models in robotics. They are commonly used as expressive multimodal generative model priors for downstream tasks, as similarly done in image generation [45]. At task planning levels, [46] proposed DALL-E-Bot, which, given a set of objects, generates a text description from a scene image and then prompts the text-to-image generator DALL-E [47] to generate a “proper goal scene”. [48] proposes StructDiffusion for arranging objects based on language commands by predicting the object arrangements using a language-conditioned diffusion model. For grasping, [49] used denoising score matching to learn a generative grasp pose model for a parallel gripper in SE(3) to optimize motion and grasping poses jointly. Several works have also used diffusion models for generating grasps using dexterous grippers [50, 51]. In trajectory planning, [52] introduced Diffuser, a trajectory generative model used for planning in Offline RL and long-horizon tasks. This work introduced a U-Net deep learning architecture to encode trajectories, which is the basis for many follow-up works that model trajectory distributions with diffusion models. In behavior cloning, [53] introduced diffusion policy, a method that learns policies that take as input a history of visual observations and outputs a sequence of actions. The action distribution is modeled with a diffusion model. Concurrently, [54] introduced BESO, which has a similar structure to diffusion policy, except they use a different formulation of score-based models. Recent methods have been built on top of these two works to include better perception, language commands, collision avoidance (for dynamic environments), equivariance and define subgoals for an MPC controller [55, 56, 57, 58, 59, 60, 61, 62]. In these works, trajectories are typically encoded using waypoints, but in [63] the authors introduced Movement Primitive Diffusion, which instead generates the parameters of a Probabilistic Dynamic Movement Primitive (ProDMP) [64], guaranteeing smoothness in the predict trajectories. Similarly, we predict the control points of a B-spline.
@@ III Robot Motion Planning with Diffusion Models
In this section, we detail our method – Motion Planning Diffusion (MPD) – and the necessary background. An overview of the inference process is illustrated in fig. 2.
@@ III-A Optimization-based Motion Planning
[…]
@@
[…]
where the costs $C_{\text{vel}}$ and $C_{\text{acc}}$ promote short and smooth paths, and are detailed in section III-G.
As optimization can become unfeasible while trying to satisfy the constraints, it is common practice to relax and include them in the cost function as soft constraints [2, 81, 3]. As we do batch optimization, we can filter the solutions for the ones respecting the constraints. Hence, we solve
@@ III-B Motion Planning as Inference
Solving eq. 1 with gradient-based methods relies on an initial trajectory, that is iteratively improved by following the cost gradient. In some cases, we would like the solution to remain close to the initialization to keep properties such as smoothness or if the initial trajectory is derived from human demonstrations. Then, instead of a single particle, it is useful to describe a prior distribution over trajectories. Hence, the motion planning problem can be formulated as inference [82, 83]. The goal is to either sample from or maximize the posterior distribution of trajectories given the task objective
@@
is equivalent to eq. 2 regularized with the prior. Contrary to classical optimization-based motion planning, planning-as-inference has several advantages. Notably, it provides a principal way to introduce informative priors to planning problems, e.g., GPMP2 [3] utilizes a GP to encode dynamic feasibility and smoothness. Additionally, specialized methods can be employed to sample from the posterior. We use diffusion models as a prior over trajectories, and instead of computing the MAP solution, sample from the posterior to obtain a trajectory distribution. Sampling from the posterior was done previously for GPMP with Gaussian approximations using variational inference, which are inherently unimodal [84]. Instead, the diffusion framework provides an approach to obtain a multimodal distribution, which allows for a more expressive posterior and a natural way to treat stochasticity.
@@ III-D Blending Sampling and Optimization
First, the entries in ${\bm{\Sigma}}_{i}$ approach 00 with $i\to 0$. Hence, the guidance function vanishes towards the end of the denoising process. This is desirable if the prior covers a very large space of the data manifold where the objective likelihood is high. However, if there is no (or little) data in those regions, the generative model might be unable to move samples towards high objective likelihood regions. To counteract this and keep the influence of the task likelihood, we drop the covariance ${\bm{\Sigma}}_{i}$. This is equivalent to changing the cost weights per denoising step $\lambda_{j}\coloneqq{\bm{\Sigma}}_{i}^{-1}\lambda_{j}$, which keeps the likelihood relevancy.
@@ III-E Trajectory Parametrization
Figure 3: The figures show the resulting trajectories when using linear phase-time scaling. In LABEL:sub@fig:phase_time_control_points, the control points (blue dots) and path of a clamped $5$th order B-spline are displayed. LABEL:sub@fig:phase_time_q_dq_ddq shows the trajectory in time of positions, velocities, and accelerations for the two degrees of freedom.
@@
To accelerate computations, similarly like in [93], we discretize the phase variable into $n_{s}$ steps and pre-compute a matrix ${{\bm{B}}\in\mathbb{R}^{n_{s}\times n_{b}}}$ of B-splines basis, where $n_{s}$ is equivalent to $H$ in the dense waypoint representation. A joint position trajectory in phase-space is given by ${\bm{Q}}={\bm{B}}{\bm{w}}\in\mathbb{R}^{n_{s}\times d}$.
@@
$2\displaystyle\dot{{\bm{q}}}(t)=\partialderivative{{\bm{q}}(s)}{s}\frac{1}{T},\quad\ddot{{\bm{q}}}(t)=\partialderivative[2]{{\bm{q}}(s)}{s}\frac{1}{T^{2}}.$
@@ III-F Implementation Details
[52] proposed a temporal U-Net architecture with $1$-dimensional convolutions to encode local relations between states and actions, acting as a low-pass filter that prevents jumps between neighboring states. Due to its success in modeling trajectories, we adapt it and replace states with B-spline control points and extend this network architecture to include a conditioning variable, along with the diffusion step index, using Feature-wise Linear Modulation (FiLM) [100] (see fig. 2). Additionally, all data is normalized to $[-1,1]$.
For sampling, we tested both DDPM and DDIM algorithms. In DDIM, the number of steps $|M|$ is commonly selected linearly from $\{1,\ldots,N\}$. With this schedule, several steps closer to $N$ add too much noise to samples during the denoising process. We can skip those first steps by choosing quadratically distributed steps, which means the denoising process jumps from larger indexes to ones closer to 00, leading to fewer denoising steps. Moreover, we use the deterministic version of DDIM with $\gamma=0$, which slightly prevents the risk of adding noise to a trajectory that is already collision-free. During inference, cost guidance is computed w.r.t. the $\bm{\tau}_{i}$ at denoising step $i$. However, for higher noise levels, the trajectory is still quite noisy, and this gradient produces little effect. Hence, we only apply the cost gradients on the last $i_{\text{cost}}$ steps of the denoising process.
@@ III-G Motion Planning Costs
[…]
where the integral is approximated by discretizing the phase variable into $n_{s}$ segments (section III-E). We consider a fixed time duration $T$ and leave its optimization for future work.
@@
[…]
@@
[…]
@@
Collision costs. The task-space signed distance function of an environment $\textsc{sdf}({\bm{x}})$ is the smallest signed Euclidean distance between a point in space ${\bm{x}}\in\mathbb{R}^{3}$ and the closest surface (negative if inside an obstacle, and positive otherwise).
@@
Given an environment with obstacles, we precompute and store the sdf using a fine voxel grid representation and project a continuous point ${\bm{x}}$ to the nearest point in the grid. The sdf representation is a common assumption in motion planning [2, 3], which can be obtained fast in the real world using newer software modules [104]. Also, if a new object is added to the scene, and we know its sdf, the resulting sdf is the minimum between the environment and the new object sdf. Similarly, we can precompute the sdf gradient $\nabla_{{\bm{x}}}\textsc{sdf}({\bm{x}})$, which has L2222-norm equal to $1$. This gradient indicates a direction to push the robot out of collision when the $\textsc{sdf}({\bm{x}})<0$. To compute robot collisions, we represent it with $S$ spheres along its body $\mathcal{B}$ (see fig. 4), allowing for faster computation of the sdf between the robot and the environment compared to using a full mesh. We consider two types of collisions: with the environment and self collisions. The environment collision cost is defined as
@@
Joint limits. Joint limits in position, velocity, and acceleration are enforced by computing the L2222-norm joint violations, with a small margin ${\epsilon\geq 0}$ on the $d$ degrees-of-freedom
@@
$\qquad\begin{cases}\frac{1}{2}\|\bm{\tau}_{d,\text{min}}(s)+\epsilon-\bm{\tau}_{d}(s)\|_{2}^{2}&\text{if }\bm{\tau}_{d}(s)<\bm{\tau}_{d,\text{min}}(s)+\epsilon\\ \frac{1}{2}\|\bm{\tau}_{d,\textrm{max}}(s)-\epsilon-\bm{\tau}_{d}(s)\|_{2}^{2}&\text{if }\bm{\tau}_{d}(s)>\bm{\tau}_{d,\text{max}}(s)-\epsilon\\ 0&\text{else}\\ \end{cases}$

### v3

@@ Abstract
The performance of optimization-based robot motion planning algorithms is highly dependent on the initial solutions, commonly obtained by running a sampling-based planner to obtain a collision-free path. However, these methods can be slow in high-dimensional and complex scenes and produce non-smooth solutions. Given previously solved path-planning problems, it is highly desirable to learn their distribution and use it as a prior for new similar problems. Several works propose utilizing this prior to bootstrap the motion planning problem, either by sampling initial solutions from it, or using its distribution in a maximum-a-posterior formulation for trajectory optimization. In this work, we introduce Motion Planning Diffusion (MPD), an algorithm that learns trajectory distribution priors with diffusion models. These generative models have shown increasing success in encoding multimodal data and have desirable properties for gradient-based motion planning, such as cost guidance. Given a motion planning problem, we construct a cost function and sample from the posterior distribution using the learned prior combined with the cost function gradients during the denoising process. Instead of learning the prior on all trajectory waypoints, we propose learning a lower-dimensional representation of a trajectory using linear motion primitives, particularly B-spline curves. This parametrization guarantees that the generated trajectory is smooth, can be interpolated at higher frequencies, and needs fewer parameters than a dense waypoint representation. We demonstrate the results of our method ranging from simple 22D to more complex tasks using a $7$-dof robot arm manipulator. In addition to learning from simulated data, we also use human demonstrations on a real-world pick-and-place task. The experiment results show that diffusion models are strong priors for encoding multimodal trajectory distributions for optimization-based motion planning. https://sites.google.com/view/motionplanningdiffusion
@@ I Introduction
The key contributions of our work are:
  • Motion Planning Diffusion (MPD) - a novel method that combines learning and adaptation of robot trajectories using diffusion models; implemented as a diffusion-based generative framework that learns trajectory priors in a compact B-spline parameterization.
  • A cost-guided posterior sampling, following the planning-as-inference framework, that blends learned trajectory priors with task-specific objectives during the reverse diffusion process, enabling the generation of diverse, feasible, and collision-free trajectories.
  • An empirical evaluation demonstrating MPD’s effectiveness across a variety of planning problems, ranging from simple 2D setups to environments with a 7-dof robot manipulator, and a real-world pick-and-place task learned from human demonstrations via kinesthetic teaching.
@@ II-A Path and motion planning
On the other hand, optimization-based motion planners search for a collision-free movement by locally optimizing a trajectory either via gradient descent or stochastic optimization, aiming for smoothness while satisfying other objective constraints. These methods can be seen as performing sampling-based planners’ search and smoothening steps in one step. Gradient-based methods include CHOMP [2], TrajOpt [4] and GPMP/GPMP2 [3]. Stochastic optimization methods based on path integral formulations [32] include STOMP [33], Stochastic GPMP [34] and MGPTO [35]. Other works solve motion planning by optimal transport [36].
@@ II-C Diffusion models in robotics and motion planning
In recent years, several works have explored score-based and diffusion models in robotics. They are commonly used as expressive multimodal generative model priors for downstream tasks, as similarly done in image generation [45]. At task planning levels, [46] proposed DALL-E-Bot, which, given a set of objects, generates a text description from a scene image and then prompts the text-to-image generator DALL-E [47] to generate a “proper goal scene”. [48] proposes StructDiffusion for arranging objects based on language commands by predicting the object arrangements using a language-conditioned diffusion model. For grasping, [49] used denoising score matching to learn a generative grasp pose model for a parallel gripper in SE(3) to optimize motion and grasping poses jointly. Several works have also used diffusion models for generating grasps using dexterous grippers [50, 51]. In trajectory planning, [52] introduced Diffuser, a trajectory generative model used for planning in Offline RL and long-horizon tasks. This work introduced a U-Net deep learning architecture to encode trajectories, which is the basis for many follow-up works that model trajectory distributions with diffusion models. MPD builds on the ideas of Diffuser but with a focus on robot motion planning. Namely, we use a B-spline trajectory representation instead of waypoints, leading to faster inference and smooth trajectories; and at inference time, we add new obstacles not seen during training and generate collision-free trajectories using the environment’s signed distance function as a cost function. In behavior cloning and visuomotor imitation learning from human demonstrations, [53] introduced diffusion policy, a method that learns policies that take as input a history of visual observations and outputs a sequence of actions. The action distribution is modeled with a diffusion model. Concurrently, [54] introduced BESO, which has a similar structure to diffusion policy, except they use a different formulation of score-based models. These works focused on closed-loop reactive policies that generate short-horizon trajectories, while we focus on generating open-loop point-to-point collision-free trajectories. Recent methods have been built on top of these two works to include better perception, language commands, collision avoidance (for dynamic environments), equivariance and define subgoals for an MPC controller [55, 56, 57, 58, 59, 60, 61, 62]. In these works, trajectories are typically encoded using waypoints, but in [63] the authors introduced Movement Primitive Diffusion, which instead generates the parameters of a Probabilistic Dynamic Movement Primitive (ProDMP) [64], guaranteeing smoothness in the predict trajectories. Similarly, we predict the control points of a B-spline.
@@ III Robot Motion Planning with Diffusion Models
This section details our method and the necessary background. Figure 2 shows an overview of the inference process.
@@ III-A Optimization-based Motion Planning
Starting from an initial joint position ${\bm{q}}_{\text{start}}\in\mathbb{R}^{d}$ with zero velocity and acceleration, we consider the task of generating a smooth joint trajectory that either reaches a goal joint position ${\bm{q}}_{\text{goal}}\in\mathbb{R}^{d}$ or a goal end-effector pose $^{W}{\bm{H}}^{EE}_{\text{goal}}\in\textsc{SE(3)}$, with zero velocity and acceleration. The obtained trajectory must avoid collisions, be smooth, short, and satisfy joint limits. Let ${\bm{q}}(0)$ and ${\bm{q}}(T)\in{\mathbb{R}}^{d}$ be the start and final joint positions of a robot with $d$ degrees-of-freedom, respectively, $T$ a fixed trajectory duration, $\dot{{\bm{q}}}$ the velocities, $\ddot{{\bm{q}}}$ the accelerations and $\bm{\tau}(t)=({\bm{q}}(t),\dot{{\bm{q}}}(t),\ddot{{\bm{q}}}(t))$} a trajectory. This problem is formulated as trajectory optimization:
@@
$\quad\text{no collisions},$
where the costs $C_{\text{vel}}$ and $C_{\text{acc}}$ promote short and smooth paths, and are detailed in section III-G. In this work we consider a fixed trajectory duration $T$, similarly to other motion planning algorithms [2, 33, 3]. The optimization of minimum-time trajectories is left for future work.
As the optimization can become unfeasible while trying to satisfy the constraints, it is common practice to relax and include them in the cost function as soft constraints [2, 81, 3]. As we do batch optimization, we can filter the solutions for the ones respecting the constraints. Hence, we solve
@@ III-B Motion Planning as Inference
Solving eq. 1 with gradient-based methods relies on an initial trajectory. In some cases, we would like the solution to remain close to the initialization to keep properties such as smoothness or if the initial trajectory is derived from human demonstrations. Then, instead of a single particle, it is useful to describe a prior distribution over trajectories. Hence, the motion planning problem can be formulated as inference [82, 83]. The goal is to either sample from or maximize the posterior distribution of trajectories given the task objective
@@
is equivalent to eq. 2 regularized with the prior. Contrary to classical optimization-based motion planning, planning-as-inference has several advantages. Notably, it provides a principled way to introduce informative priors to planning problems, e.g., GPMP2 [3] utilizes a GP to encode dynamic feasibility and smoothness. Additionally, specialized methods can be employed to sample from the posterior. We use diffusion models as a prior over trajectories, and instead of computing the MAP solution, sample from the posterior to obtain a trajectory distribution. Sampling from the posterior was done previously for GPMP with Gaussian approximations using variational inference, which are inherently unimodal [84]. Instead, the diffusion framework provides an approach to obtain a multimodal distribution, which allows for a more expressive posterior and a natural way to treat stochasticity.
@@ III-D Blending Sampling and Optimization
First, the entries in ${\bm{\Sigma}}_{i}$ approach 0 with $i\to 0$. Hence, the guidance function vanishes towards the end of the denoising process. This is desirable if the prior covers a very large space of the data manifold where the objective likelihood is high. However, if there is no (or little) data in those regions, the generative model might be unable to move samples towards high objective likelihood regions. To counteract this and keep the influence of the task likelihood, we drop the covariance ${\bm{\Sigma}}_{i}$. This is equivalent to changing the cost weights per denoising step $\lambda_{j}\coloneqq{\bm{\Sigma}}_{i}^{-1}\lambda_{j}$, which keeps the likelihood relevancy.
@@ III-E Trajectory Parametrization
Figure 3: The figures show the resulting trajectories when using linear phase-time scaling. In LABEL:sub@fig:phase_time_control_points, the control points (blue dots) and path of a clamped 55th order B-spline are displayed. LABEL:sub@fig:phase_time_q_dq_ddq shows the trajectory in time of positions, velocities, and accelerations for the two degrees of freedom.
@@
Note that if ${\bm{q}}_{\text{goal}}$ is not specified, but rather a desired end-effector pose $^{W}{\bm{H}}^{EE}_{\text{goal}}$, then the last control point ${\bm{w}}_{n_{b}-1}$ is generated by the diffusion model. To accelerate computations, similarly to [93], we discretize the phase variable into $n_{s}$ steps and pre-compute a matrix ${\bm{B}}\in\mathbb{R}^{n_{s}\times n_{b}}$} of B-splines basis, where $n_{s}$ is equivalent to $H$ in the dense waypoint representation. A joint position trajectory in phase-space is given by ${\bm{Q}}={\bm{B}}{\bm{w}}\in\mathbb{R}^{n_{s}\times d}$.
@@
[…]
@@ III-F Implementation Details
[52] proposed a temporal U-Net architecture with $1$-dimensional convolutions to encode local relations between states and actions, acting as a low-pass filter that prevents jumps between neighboring states. Due to its success in modeling trajectories, we adapt it and replace states with B-spline control points and extend this network architecture to include a conditioning variable, along with the diffusion step index, using Feature-wise Linear Modulation (FiLM) [100] (see fig. 2). Additionally, all data is normalized to [1,1][-1,1].
For sampling, we tested both DDPM and DDIM algorithms. In DDIM, the number of steps $|M|$ is commonly selected linearly from $\{1,\ldots,N\}$. With this schedule, several steps closer to $N$ add too much noise to samples during the denoising process. We can skip those first steps by choosing quadratically distributed steps, which means the denoising process jumps from larger indexes to ones closer to 0, leading to fewer denoising steps. Moreover, we use the deterministic version of DDIM with $\gamma=0$, which slightly prevents the risk of adding noise to a trajectory that is already collision-free. During inference, cost guidance is computed w.r.t. the $\bm{\tau}_{i}$ at denoising step $i$. However, for higher noise levels, the trajectory is still quite noisy, and this gradient produces little effect. Hence, we only apply the cost gradients on the last $i_{\text{cost}}$ steps of the denoising process.
@@ III-G Motion Planning Costs
where the integral is approximated by discretizing the phase variable into $n_{s}$ segments (section III-E).
@@
where $C_{j}\in\mathbb{R}$, $\partial C({\bm{q}})/\partial{\bm{q}}\in\mathbb{R}^{1\times d}$, and for B-splines $d\differential{\bm{q}}/\differential{\bm{w}}={\bm{B}}\in\mathbb{R}^{d\times n_{b}\times d}$, where the basis matrix is repeated along the degrees-of-freedom to match dimensions. If $C_{j}$ is a task space cost, let ${\bm{X}}_{m}=\text{FK}_{m}({\bm{q}})\in\textsc{SE(3)}$ be the pose of a link $m$ on the robot arm computed with forward kinematics, and consider a cost $C_{j}({\bm{X}}_{m})$. Then we have
@@
$d_{\textsc{SE(3)}}$ defines the distance between two elements of SE(3). Given two poses (homogeneous transformations) ${{\bm{T}}_{1}=[{\bm{R}}_{1},{\bm{p}}_{1}]}\in\textsc{SE(3)}$ and ${{\bm{T}}_{2}=[{\bm{R}}_{2},{\bm{p}}_{2}]}\in\textsc{SE(3)}$, consisting of a translational and rotational part, we choose
@@
Collision costs. The task-space signed distance function of an environment $\textsc{sdf}({\bm{x}})$ is the
@@
smallest signed Euclidean distance between a point in space ${\bm{x}}\in\mathbb{R}^{3}$ and the closest surface (negative if inside an obstacle, and positive otherwise). Given an environment with obstacles, we precompute and store the sdf using a fine voxel grid representation and project a continuous point $\bm{x}$} to the nearest point in the grid. The sdf representation is a common assumption in motion planning [2, 3], which can be obtained fast in the real world using newer software modules [104]. Also, if a new object is added to the scene, and we know its sdf, the resulting sdf is the minimum between the environment and the new object sdf. Similarly, we can precompute the sdf gradient $\nabla_{{\bm{x}}}\textsc{sdf}({\bm{x}})$, which has L22-norm equal to $1$. This gradient indicates a direction to push the robot out of collision when the $\textsc{sdf}({\bm{x}})<0$. To compute robot collisions, we represent it with $S$ spheres along its body $\mathcal{B}$ (see fig. 4), allowing for faster computation of the sdf between the robot and the environment compared to using a full mesh. We consider two types of collisions: with the environment and self collisions. The environment collision cost is defined as
@@
Joint limits. Joint limits in position, velocity, and acceleration are enforced by computing the L22-norm joint violations, with a small margin $\epsilon\geq 0$} on the $d$ degrees-of-freedom
@@
[…]